A deuteron, 21H, is the nucleus of a hydrogen isotope and consists of one proton and one neutron. The plasma of deuterons in a nuclear fusion reactor must be heated to about 3.10×108 K .

A) What is the rms speed of the deuterons?

B) Is this a significant fraction of the speed of light ( c = 3.0×108m/s )? Find the percentage! (Vrms/C)
Answer #1

A) vrms = sqrt(3RT/M)

where R = universal gas constant

T = temperature in kelvin

M = molecular mass = 2 g for deutron

V = sqrt(3*8.314*3.1*10^8 / 2*10^-3)

= 1.96*10^6 m/s

B) 1.96*10^6 / 3*10^8 = 0.0065

= 0.655 % of the speed of light

so this is not a significant fraction